OpenStudy (anonymous):

Check my answer for me please? Mrs. Miller sells a house for $179,000. If she earns a commission of 6%, how much money does she earn? Write a proportion and show your work.

OpenStudy (anonymous):

\[\frac{ x }{ 179,000 } = \frac{ 6 }{ 100 }\] \[\frac{ 100x }{ 100 } = \frac{ 1,074,000 }{ 100 }\] \[x = 10,740\]

OpenStudy (anonymous):

thats right :D

OpenStudy (anonymous):

Thank you! @ash.bug